# Dynamical Constraints on the Vertical Structure of Jupiter’s Polar Cyclones

### v1

@@ Acknowledgments
$f=2\Omega\cos\left(\frac{\sqrt{x^{2}+y^{2}}}{R_{\rm{J}}}\right),$ (S16)
$\Omega=1.759\times 10^{-4}$ is Jupiter’s rotation rate, and $R_{\mathrm{J}}=66{,}854$ km is Jupiter’s radius. In practice, our model solves the non-dimensional form of Eq. S15:
$\frac{D}{Dt}\left(\nabla^{2}\psi+\hat{f}-\frac{1}{\rm Bu}\psi\right)=\rm{Ek}\nabla^{4}\psi,$ (S17)
where lengths are scaled by $L=1{,}000$ km and velocities by $U=50$ m s-1. Thus, time is scaled by $T=L/U$. The non-dimensional variables are $\hat{\psi}=\tfrac{1}{LU}\,\psi$ and $\hat{f}=\tfrac{L}{U}\,G(r,R_{\mathrm{d}},A_{\mathrm{d}})\,f$. We set the numerical viscosity to $\mathrm{Ek}=10^{-5}$, and the Burger number is $\mathrm{Bu}=(L_{d}/L)^{2}$, where $L_{d}$ is an input parameter.
We solve the model using Dedalus Burns et al. (2020) on a doubly periodic Fourier domain. The initial condition consists of $N_{\mathrm{cyc}}$ cyclones (9 for the north pole and 6 for the south pole) of the form (in dimensional variables):
$\xi=\sum_{n=1}^{N_{\rm{cyc}}}\frac{V}{R}\left(2-\left(\frac{r_{n}}{R}\right)^{b}\right)\exp\left[\frac{1}{b}\left(1-\left(\frac{r_{n}}{R}\right)^{b}\right)\right],$ (S18)
where $r_{n}=\sqrt{(x-x_{n})^{2}+(y-y_{n})^{2}}$ measures the distance from the center of cyclone $n$. The cyclone centers $\{(x_{n},y_{n})\}$ are taken from Juno PJ4 observations Mura et al. (2022). We obtain the initial streamfunction by solving $\nabla^{2}\psi=\xi$ prior to each run.
To mitigate boundary effects from the rectangular domain, we use the "trap" method Siegelman et al. (2022a). Specifically, we multiply the planetary PV by an activation function (Eq. S13) which smoothly transitions from $\approx 0$ outside $r>R_{\mathrm{d}}$ to $\approx\tfrac{L}{U}f$ inside $r<R_{\mathrm{d}}$. In this study, we use $A_{\mathrm{d}}=-20$ (negative values switch the domain of activation to within the circle), and choose $R_{\mathrm{d}}$ to be 97% of the total domain size. This procedure effectively "disconnects" the dynamics of interest (inside $R_{\mathrm{d}}$) from the spurious influence of the periodic boundaries (outside $R_{\mathrm{d}}$). Throughout each simulation, we track the centers of the cyclones in time and maintain their individual identities, thereby obtaining a record of their trajectories and possible mergers.
@@
Because we sample the model repeatedly, we first identified the smallest resolution and domain size that would produce acceptably accurate cyclone trajectories. As a benchmark, we ran a high-resolution model ($360\times 360$ grid points) with a large domain ($36{,}000$ km on each side). We initialized nine identical cyclones at the north polar PJ4 positions $\bigl{(}L_{d}=348\;\mathrm{km},R=867\;\mathrm{km},V=86.13\;\mathrm{m\,s}^{-1},b=1.51\bigr{)}$, and integrated for $2{,}500$ days.
Next, we reran the model at coarser resolutions and smaller domain sizes. For each simulation, we computed the RMSE between the resulting cyclone trajectories and those from the benchmark (Fig. S1). We are primarily interested in the average westward drift of these cyclones, so we allowed a maximum mean trajectory error of 100 km (blue contour in Fig. S1a). We found that a $72\times 72$ grid with a domain size of $30{,}400$ km satisfies this requirement. Figure S1b,c illustrates that the associated trajectories agree well with the benchmark. We therefore adopted these numerical settings for the sampling runs used to generate Fig. 2.
Figure S1: Analyzing the required resolution and domain size for sampling the model. (a) Contours of RMSE (in km) between each model’s cyclone trajectories and the benchmark, as functions of resolution (abscissa) and domain size (ordinate). The blue contour highlights the 100 km threshold for acceptable error. The black dot is the chosen setting. (b) RMSE dependence on domain size for selected resolutions. (c) Benchmark trajectories (solid lines) compared with those at the chosen $72\times 72$ resolution and a $30{,}400$ km domain (dashed lines).
@@
To constrain $L_{d}$ (the deformation radius) for Jupiter’s north and south poles using the observed westward drift rates, we apply Bayesian nested sampling via the Ultranest library Buchner (2021). This algorithm explores parameter space efficiently and returns a posterior distribution of unknown parameters.
While noting that the model is idealized, and that we assume equivalent cyclones in order to drastically reduce the number of unknown variables, we are left with the 3 cyclone variables ($R,V,b$) and $L_{d}$ as our 4 unknown variables. We assume the following normal priors:
$L_{d}$ $\sim\mathcal{N}(400,150^{2})\quad\rm{(\rm{km})},$ (S19)
$b$ $\sim\mathcal{N}(1.2,0.2^{2}),$
$U$ $\sim\mathcal{N}(80,15^{2})\quad\rm{(\rm{m/s})},$
$R$ $\sim\mathcal{N}(900,150^{2})\quad\rm{(\rm{km})}.$
@@
$C_{1}=-\frac{1}{2}\left(\dfrac{\overline{\mathrm{WW}}_{\rm O}-\overline{\mathrm{WW}}_{\rm M}}{5\times 10^{-3}\,\mathrm{m\,s}^{-1}}\right)^{2}$ (S20)
@@
$C_{2}=-\frac{1}{2}\left(\dfrac{N_{\rm mergers}}{2}\right)^{2},$ (S21)
@@
and we use this total likelihood to guide the nested-sampling algorithm. Each pole (north and south) is sampled with $10{,}000$ allowed evaluations, yielding the posterior distributions presented in Fig. 2k-n.

### v2

@@ Appendix B The Polar QG Model
$f=2\Omega\cos\left(\frac{\sqrt{x^{2}+y^{2}}}{R_{\rm{Jo}}}\right),$ (24)
and $\Omega=1.759\times 10^{-4}$ is Jupiter’s rotation rate. While Eq. 24 assumes a spherical planet, Jupiter’s oblateness necessitates a correction. We incorporate this by using the osculating radius, $R_{\rm Jo}=R_{\rm Je}^{2}/R_{\rm Jp}=76{,}452$ km, where $R_{\rm Je}=71{,}492$ km and $R_{\rm Jp}=66{,}854$ km are the equatorial and polar radii, respectively Hyder et al. (2022). This adjustment ensures the Coriolis term reflects the correct curvature near the polar regions. We note that, for translating between latitude–longitude and Cartesian (xxyy) coordinates of the cyclones, we use Jupiter’s polar radius, which is more appropriate given the local geometry near the poles.
In practice, our model solves the non-dimensional form of Eq. 3:
$\frac{D}{Dt}\left(\nabla^{2}\psi+\hat{f}-\frac{1}{\rm Bu}\psi\right)=\rm{Ek}\nabla^{4}\psi,$ (25)
where lengths are scaled by L=1, [^0] L=1{,}000 km and velocities by $U=50$ m s-1. Thus, time is scaled by $T=L/U$. The non-dimensional variables are $\hat{\psi}=\tfrac{1}{LU}\,\psi$ and $\hat{f}=\tfrac{L}{U}\,G(r,R_{\mathrm{d}},A_{\mathrm{d}})\,f$. We set the numerical viscosity to $\mathrm{Ek}=10^{-5}$, and the Burger number is $\mathrm{Bu}=(L_{d}/L)^{2}$, where $L_{d}$ is an input parameter.
We solve the model using Dedalus Burns et al. (2020) on a doubly periodic Fourier domain. The initial conditions consist of $N_{\mathrm{cyc}}$ cyclones (9 for the north pole and 6 for the south pole) of the form (in dimensional variables):
$\xi_{(t=0)}=\sum_{n=1}^{N_{\rm{cyc}}}\frac{V}{R}\left(2-\left(\frac{r_{n}}{R}\right)^{b}\right)\exp\left[\frac{1}{b}\left(1-\left(\frac{r_{n}}{R}\right)^{b}\right)\right],$ (26)
where $r_{n}=\sqrt{(x-x_{n})^{2}+(y-y_{n})^{2}}$ measures the distance from the center of cyclone $n$. The cyclone centers $\{(x_{n},y_{n})\}$ are taken from Juno PJ4 observations Mura et al. (2022), with the original planetocentric latitude coordinates converted to Cartesian positions using Jupiter’s polar radius, $R_{\rm Jp}$. We obtain the initial streamfunction by solving $\nabla^{2}\psi=\xi_{(t=0)}$ prior to each run.
To mitigate boundary effects from the rectangular domain, we use the "trap" method Siegelman et al. (2022a). Specifically, we multiply the planetary PV by an activation function (Eq. 21) which smoothly transitions from $\approx 0$ outside $r>R_{\mathrm{d}}$ to $\approx\tfrac{L}{U}f$ inside $r<R_{\mathrm{d}}$. In this study, we use $A_{\mathrm{d}}=-20$ (negative values switch the domain of activation to within the circle), and choose $R_{\mathrm{d}}$ to be 97% of the total domain size. This procedure effectively "disconnects" the dynamics of interest (inside $R_{\mathrm{d}}$) from the spurious influence of the periodic boundaries (outside $R_{\mathrm{d}}$). Throughout each simulation, we track the centers of the cyclones in time and maintain their individual identities, thereby obtaining a record of their trajectories and possible mergers. A Dedalus model script for solving Eqs. 2526 is available on GitHub (polar_cyclones.py) along with the tracking analysis script (polar_cyclones_cost.py).
@@
Because we sample the model repeatedly, we first identified the smallest resolution and domain size that would produce acceptably accurate cyclone trajectories. As a benchmark, we ran a high-resolution model ($360\times 360$ grid points) with a large domain (36,00036{,}000 km on each side). We initialized nine identical cyclones at the north polar PJ4 positions $\bigl(L_{d}=348\;\mathrm{km},R=867\;\mathrm{km},V=86.13\;\mathrm{m\,s}^{-1},b=1.51\bigr)$, and integrated for 2,5002{,}500 days.
Next, we reran the model at coarser resolutions and smaller domain sizes. For each simulation, we computed the RMSE between the resulting cyclone trajectories and those from the benchmark (Fig. S3). We are primarily interested in the average westward drift of these cyclones, so we allowed a maximum mean trajectory error of 100 km (blue contour in Fig. S3a). We found that a $96\times 96$ grid with a domain size of 30,40030{,}400 km satisfies this requirement. Figure S3b,c illustrates that the associated trajectories agree well with the benchmark. We therefore adopted these numerical settings for the sampling runs used to generate Fig. 2.
Figure S3: Analyzing the required resolution and domain size for sampling the model. (a) Contours of RMSE (in km) between each model’s cyclone trajectories and the benchmark, as functions of resolution (abscissa) and domain size (ordinate). The blue contour highlights the 100 km threshold for acceptable error. The black dot is the chosen setting. (b) RMSE dependence on domain size for selected resolutions. (c) Benchmark trajectories (solid lines) compared with those at the chosen $96\times 96$ resolution and a 30,40030{,}400 km domain (dashed lines).
@@
To constrain $L_{d}$ (the deformation radius) for Jupiter’s north and south poles using the observed westward drift rates, we apply Bayesian nested sampling via the Ultranest library Buchner (2021). This algorithm explores parameter space efficiently, according to a likelihood function, and returns a posterior distribution of unknown parameters.
While noting that the model is idealized, and that we assume equivalent cyclones in order to drastically reduce the number of unknown variables, we are left with the 3 cyclone variables ($R,V,b$) and $L_{d}$ as our 4 unknown variables. We assign the following normal prior distributions to these parameters, denoted $\mathcal{N}(\mu,\sigma^{2})$, where $\mu$ is the mean and $\sigma$ is the standard deviation of the distribution:
[…]
@@
$C_{1}=-\frac{1}{2}\left(\dfrac{\overline{\mathrm{WW}}_{\rm O}-\overline{\mathrm{WW}}_{\rm M}}{5\times 10^{-3}\,\mathrm{m\,s}^{-1}}\right)^{2}$ (28)
@@
$C_{2}=-\frac{1}{2}\left(\dfrac{N_{\rm mergers}}{2}\right)^{2},$ (29)
@@
and we use this total likelihood to guide the nested-sampling algorithm. Each pole (north and south) is sampled with 15,00015{,}000 allowed evaluations, yielding the posterior distributions presented in Fig. 2k-n. The joint posterior distributions, shown as corner plots in Fig. S4, reveal that among all parameter pairs, only $b$ and $R$ exhibit a simple linear correlation, while the remaining parameters display more complex relationships.
Figure S4: Joint and marginal posterior distributions of the model parameters for the North Pole (blue) and South Pole (orange). Diagonal panels are the same as Fig. 2k-n; off-diagonal panels display the joint densities with filled contours at the $10\%$, $68.27\%\,(1\sigma)$ and $95.45\%\,(2\sigma)$ levels overlaid with individual samples (points).
Appendix C The Brunt-Väisälä Frequency of Jupiter’s Atmosphere
Figure S5: The Brunt Väisälä (NN) frequency from the Galileo probe decent. (a) The static stability, $S$, digitized from Fig. 7 in (Magalhães, 2002). (b) Calculation of NN based on panel (a). The red vertical line indicates the pressure-weighted mean value of NN over the range $1\leq p\leq 22$ bar.
The results shown in Fig. 3 depend on the vertical profile of the Brunt-Väisälä frequency ($N$, in units of s-1). The only direct measurement of temperature at pressures greater than 1 bar comes from the Galileo probe, which entered Jupiter’s atmosphere on December 7, 1995, at $6.5^{\circ}$N and transmitted data down to approximately 22 bar (Seiff et al., 1996). A detailed analysis of the vertical static stability from this measurement was presented by Magalhães et al.(2002) (Magalhães, 2002), and is digitized in Fig. S5a.
In their analysis, static stability is expressed as the difference between the vertical temperature gradient and the dry adiabatic lapse rate:
$S=\frac{\partial T}{\partial z}+\Gamma\left(\mathrm{K}\;\mathrm{km}^{-1}\right).$ (31)
We then compute $N$ (Fig. S5b) using the relation:
$N^{2}=\frac{g}{T}\left(\frac{\partial T}{\partial z}+\Gamma\right)=\frac{g}{T}S.$ (32)
However, there are several caveats to using this profile in our calculations (Fig. 3). First, the Galileo probe descended at an equatorial latitude ($6.5^{\circ}$N), which is dynamically distinct from the polar region of interest. Second, the probe entered a known atmospheric “hot spot” (Magalhães, 2002), which may not be representative of Jupiter’s mean atmospheric structure. Third, significant uncertainties affect the measurement itself, including unknown vertical wind velocities during descent and thermal disturbances near the probe’s sensors. These factors may explain the (likely spurious) negative static stability values observed near $10$-bar.
Given these uncertainties, the limited depth range of the data, and to ensure consistency and simplicity in our eigenvalue calculations (Fig. 3), we opt to use a constant value of $N$. Nevertheless, the pressure-weighted average,
$\langle N\rangle=\frac{\int_{1\rm{bar}}^{22\penalty 10000\ \rm{bar}}N(p)\,p\,dp}{\int_{1\penalty 10000\ \rm{bar}}^{22\penalty 10000\ \rm{bar}}p\,dp}\sim 3.7\times 10^{-3}\mathrm{s}^{-1},$ (33)
agrees well with estimates by Lee et al.(2021) (Lee and Kaspi, 2021), inferred from Juno-based ammonia distribution, and is referenced accordingly in this manuscript.
Appendix D Equivalence Between the 3D and 2D QG Models for a Given Mode
In this section, we provide a mathematical argument for the equivalence between the 2D-QG model results shown in Fig. 2 and the 3D-QG model decomposed into vertical modes, as presented in Fig. 3.
We begin with the 3D-QG equation Vallis (2017):
$\frac{D_{h}}{Dt}\left(\nabla_{h}^{2}\psi+f+\frac{f_{0}^{2}}{\tilde{\rho}}\frac{\partial}{\partial z}\left(\frac{\tilde{\rho}}{N^{2}}\frac{\partial\psi}{\partial z}\right)\right)=0,$ (34)
where the subscript ’$h$’ denotes derivatives taken only in the horizontal directions.
Assuming separability of the streamfunction into horizontal and vertical components, we write,
$\psi(x,y,z,t)=\sum_{n}\psi_{n}(x,y,t)\Phi_{n}(z),$ (35)
where the vertical modes $\Phi_{n}(z)$ are required to be orthogonal:
$\langle\Phi_{n},\Phi_{m}\rangle\equiv\int_{-H}^{0}w\,\Phi_{n}\Phi_{m}\;dz=C_{nm}\delta_{nm},$ (36)
with $w(z)=\tilde{\rho}(z)$ serving as a weight function, $C_{nm}$ constant coefficients, and $\delta_{nm}$ the Kronecker delta. Such orthogonal modes are given by $\Phi_{n}$, which are non-trivial solutions of the Sturm-Liouville equation,
subject to the boundary conditions:
$\partial_{z}\Phi_{n}(0)=0,\qquad\Phi_{n}(-H)=0.$ (38)
These solutions are thus eigenfunctions ($\Phi_{n}$ for mode $n$) of Eqs. 37-38, where the eigenvalues are $\Gamma_{n}$.
Substituting the modal decomposition (Eqs. 3537) into Eq. 34 yields
$\frac{D_{h}}{Dt}\left(\sum_{n}\Phi_{n}\left(\nabla^{2}_{h}\psi_{n}-\Gamma_{n}\psi_{n}\right)+f\right)=0.$ (39)
Linearizing Eq. 39, and projecting onto a single mode by taking the inner product, $\Phi_{n}$, Eq. 39>, gives
$\frac{\partial}{\partial t}\left(\nabla^{2}\psi_{n}-\Gamma_{n}\psi_{n}\right)+J(\psi_{n},f)=0,$ (40)
where $J(a,b)=a_{x}b_{y}-a_{y}b_{x}$ denotes the Jacobian.
Thus, in the linear regime, each vertical mode of the 3D-QG system evolves independently according to a two-dimensional QG-like equation. This structure mirrors the dynamics described by the linearized 2D-QG equation:
$\frac{\partial}{\partial t}\left(\nabla^{2}\psi-\frac{1}{L_{d}^{2}}\psi\right)+J(\psi,f)=0,$ (41)
provided that $L_{d}=\Gamma_{n}^{-1/2}$. Therefore, for each vertical mode $n$, the linearized 3D-QG dynamics (Eq. 34) projected onto $\Phi_{n}(z)$ are mathematically equivalent to the 2D-QG dynamics (Eq. 41) with a mode-dependent deformation radius $L_{d}$. We argue that, while nonlinear interactions and the coexistence of multiple modes could lead to deviations from this simple structure, the leading-order dependence on $N$ and $L_{d}$ captured by this modal decomposition provides a robust framework for interpreting the vertical structures expected to be observed by Juno, as shown in Fig. 3.
Including Non-Linear Terms: Triple Interactions
Performing a modal projection, $\Phi_{n}$, Eq. 39>, without linearization, yields
$K+\frac{\partial}{\partial t}\bigl(\nabla^{2}\psi_{n}-\Gamma_{n}\psi_{n}\bigr)+J(\psi_{n},f)=0,$ (42)
where
$K=\sum_{i}\sum_{j}\Bigr(\varepsilon_{ijn}J\left({\psi_{i}},\left({\nabla{\psi_{j}}}+\Gamma_{j}\psi_{j}\right)\right)\Bigr),$ (43)
contains both self-modal and inter-modal advection terms,
$\varepsilon_{ijn}=\frac{1}{C_{nn}}\int_{-H}^{0}\Phi_{i}\Phi_{j}\Phi_{n}\,\tilde{\rho}\,dz,$ (44)
is the triple-interaction coefficient (Smith and Vallis, 2001), and the integral is symmetric under permutations of $(i,j,n)$.
For our two-mode expansion, using the eigenfunctions from Fig. 4 (with H=50.19H=50.19\;km and $N=3\times 10^{-3}\;$s-1), the nonzero elements for mode 0 are
$\varepsilon_{0ij}=\frac{1}{C_{00}}\int_{-H}^{0}\Phi_{0}\Phi_{i}\Phi_{j}\tilde{\rho}(z)dz=\left[\begin{array}[]{ccc}i\setminus j&0&1\\ 0&0.797&0.178\\ 1&0.178&0.508\end{array}\right].$ (45)
Ideally, to achieve an exact correspondence between the 2D and modal 3D QG equations, we would require
$\varepsilon_{0ij}=\begin{cases}1&\text{if }i=j=0,\\ 0&\text{otherwise}.\end{cases}$
In our case, $\varepsilon_{000}=0.797$ (quantifying the self-advection of mode 0 potential vorticity by the mode 0 velocity field), while the inter-modal coefficients are smaller.
Despite this partial modal leakage, the overall structure of the 3D QG equation in modal form retains a strong similarity to the 2D QG system, even when nonlinear interactions are included.